Components of buffer solution of 0.1 M HCn and 0.2 M Nacn. What is Ph of solution.

Grade:12

2 Answers

Arun
6 years ago
The chemical equation for this reaction is

HF + NaCN = NaF + HCN

HCN is much more weak acid than HF because it's pKa is greater

If the volumes of HF and NaCN were equal, HF is limiting reagent and is completely transformed into NaF. In the solution there is buffer HCN/CN-, and the pH of this buffer is

pH = pKa + lg[A-]/[HA]
pH = 9.4 + lg([CN-]/([HCN]) = 9.4 + lg (0.1/0.1) = 9.4
Answer: pH = 9.4
Chhavi Jain
6 years ago
HCN is a weak acid and CN- is its conjugate base . Therefore , it is an acidic buffer in which HCN is a weak acid and NaCN is providing its conjugate base (CN-).
 
To calculate the pH of an acidic buffer we will use the formula  :
 
pH= pKa + log[conjugate base]/ [acid]
 
pH= pKa + log[ CN-]/ [HCN]
 
using the data pKa of HCN at 25o C  = 9.2 ( not mentioned in the question )
pH= 9.2 + log [0.2]/ [0.1]
 
pH= 9.2+ log 2
pH= 9.2 + 0.3010 = 9.501 = 9.5 ( approx )